# Unlocking Aha Moments via Reinforcement Learning: Advancing Collaborative Visual Comprehension and Generation

### v1

@@ Abstract
11footnotetext:  Equal Contribution. 22footnotetext:  Corresponding Author.
@@ 2 Related Work
Recent studies focus on unifying visual comprehension and generation within a unified MLLM. Some efforts ge2024seed ; sun2024generative ; tong2024metamorph cascade external diffusion models after the output of MLLMs for visual generation, while some approaches chen2025janus ; pan2025generative ; zhuang2025vargpt ; wang2024emu3 tokenize images into a discrete space and then conducte unified AR for text and vision. Others xie2024show ; zhou2024transfusion try to integrate the objective of AR and diffusion into a single model. However, visual comprehension and generation still remain largely independent without collaborative synergy, and the strong reasoning mechanism of LLMs fails to revolutionize visual generation. In this paper, we aim to promote collaboration between these two capabilities, enabling a genuine CoT for visual generation and unlocking its Aha moments.
@@ 3.1 Supervised Fine-Tuning
Task-II: Self-evaluation of Text-Image Consistency: We aim to enhance the capability of MLLM to determine whether given images are semantically consistent with the text, also providing the rationale for its judgment. To construct training data, for each prompt we first select $(t-1)$ images $\{\mathcal{I}^{j},\mathcal{R}^{j}\}_{j=1}^{t-1}$ with $\mathcal{S}<0.5$ as the preceding context ($t\geq 1$). Subsequently, we randomly select one positive image ($\mathcal{S}\geq 0.7$) and one negative image ($\mathcal{S}<0.5$), respectively, as the target image $\mathcal{I}_{t}$ to be evaluated. The objective function is $p(y^{\mathcal{T}^{t}})=\sum_{j=1}^{S^{\mathcal{T}^{t}}}\log P_{\theta}(y^{\mathcal{T}^{t}}_{j}|y^{\mathcal{T}^{t}}_{<j},\mathcal{P},\{\mathcal{I}^{k},\texttt{No},\mathcal{R}^{k}\}_{k=1}^{t-1},\mathcal{I}^{t-1})$, where $\mathcal{T}^{t}$ is $(\text{Yes},\mathcal{R}^{t})$ for positive images and $(\text{No},\mathcal{R}^{t})$ for negative images.
@@ 3.2 Reinforcement Learning
(2) The comprehension reward $\texttt{R}^{Comp}=\sum_{i=1}^{K}\big{(}\mathbf{1}-|\texttt{R}^{QA}(\mathcal{I}^{i})-\texttt{SE}(\mathcal{I}^{i})|\big{)}\times T/K$, where SE()=0or $\texttt{SE}(\mathcal{I})=0\ \text{or}\ 1$ is the self-assessment on whether the generated image is semantically aligned with the text.
@@
$\rho_{i,j}$ $=\frac{\pi_{\theta}(y_{i,j}\mid\mathcal{P},y_{i,<j})}{\pi_{\theta_{\text{old}}}(y_{i,j}\mid\mathcal{P},y_{i,<j})}=\left\{\begin{array}[]{ll}\frac{\pi_{\theta}\big{(}y^{\mathcal{I}^{t}}_{i,j}|y^{\mathcal{I}^{t}}_{i,<j},\mathcal{P},\{\mathcal{I}^{k}_{i},\texttt{No},\mathcal{R}^{k}_{i}\}_{k=1}^{t-1}\big{)}}{\pi_{\theta_{\text{old}}}\big{(}y^{\mathcal{I}^{t}}_{i,j}|y^{\mathcal{I}^{t}}_{i,<j},\mathcal{P},\{\mathcal{I}^{k}_{i},\texttt{No},\mathcal{R}^{k}_{i}\}_{k=1}^{t-1}\big{)}},&\text{output the image in $t$-th round}\\[10.0pt] \frac{\pi_{\theta}(y^{\mathcal{T}^{t}}_{i,j}|y^{\mathcal{T}^{t}}_{i,<j},\mathcal{P},\{\mathcal{I}^{k}_{i},\texttt{No},\mathcal{R}^{k}_{i}\}_{k=1}^{t-1},\mathcal{I}^{t}_{i})}{\pi_{\theta_{\text{old}}}(y^{\mathcal{T}^{t}}_{i,j}|y^{\mathcal{T}^{t}}_{i,<j},\mathcal{P},\{\mathcal{I}^{k}_{i},\texttt{No},\mathcal{R}^{k}_{i}\}_{k=1}^{t-1},\mathcal{I}^{t}_{i})},&\text{self-check the image in $t$-th round}\\ \end{array}\right.$ (3)
@@ 4 Experiments
We employ Janus-Pro-7B as the backbone, developing Janus-Pro-R1 and Janus-Pro-R1-Edit, excelling in text-to-image generation (§ 4.1), image editing (§ 4.2), and image semantic evaluation (§ 4.3). More details are given in Appendix C and  D.
@@ 4.1 Introspective Text-to-Image Generation
To evaluate the capabilities of introspective image synthesis, we first conduct an automated metric evaluation on 3 text-to-image benchmarks: GenEval ghosh2023geneval , T2I-CompBench huang2023t2i , and DPG-Bench hu2024ella . The comparison results against both diffusion-based and MLLM-based methods, as well as methods incorporating CoT into image generation, are presented in Table 1. We have the following observations:
@@
In Figure 4, we present qualitative examples of Janus-Pro-R1 to trigger Aha moments within its reasoning chains to generate superior images. The model could leverage its visual comprehension capabilities to accurately identify the issues in its initial-generated images, then unleash the visual generation capabilities to output a more accurate image. Even if the newly generated image still fails to meet the requirements, the model can trigger a second Aha moment, re-evaluating the issues and repeating the visual generation to produce a fully compliant image.
Furthermore, Figure 3 presents a direct qualitative comparison between Janus-Pro-7B and our Janus-Pro-R1-7B on the final generated image for text-to-image generation tasks. It can be observed that compared to Janus-Pro-7B, after unlocking the Aha moments with CoT via a two-stage training paradigm, our model not only generates images that are more semantically aligned with the text but also achieves higher aesthetic quality.
@@
Method T2I Model Distance $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ MSE $\downarrow$ SSIM $\uparrow$ Whole $\uparrow$ Edited $\uparrow$
@@ 4.2 Image Editing
Figure 5: Qualitative examples of multi-turn editing. Our model achieves a stable trade-off between fidelity and editability.
@@ 4.3 Janus-Pro-R1 as a Stronger Image Semantic Evaluator
Furthermore, we collect several hundred pairs of semantically mismatched text-image pairs. We require the above four models to determine whether the semantics of the image and text are matched and to provide relevant reasons. We then use GPT-4o as the evaluation model to score the reasons provided by the four models from 00 (unreasonable) to $1$ (reasonable). As shown in Table 3, Janus-Pro-R1 outperforms both Janus-Pro-7B and Janus-Pro-SFT in providing reasons that are deemed more reliable by GPT-4o. It even approaches the performance of the comprehension-only MLLM, i.e., InternVL2.5-8B. These results indicate that (1) compared to the backbone model and even a comprehension-only MLLM, Janus-Pro-R1 more closely aligns with the standard metrics within GenEval, achieving greater evaluation precision; and (2) the visual comprehension capability of our model is significantly enhanced after reinforcement learning. In contrast, SFT only endows the MLLM with the basic capability for combining visual comprehension and generation, resulting in only incremental improvements to its inherent multimodal understanding ability.
@@ 4.4 In-Depth Analysis
During SFT, when selecting positive images for subtask training, we establish a higher threshold larger than 0.5 of the semantic consistent score ($\mathcal{S}\geq 0.7$ for comprehension tasks and $\mathcal{S}\geq 0.8$ for generation tasks). Though a higher threshold reduces the quantity of available training data, we find that it is crucial for improving performance. In contrast, we experiment with setting the threshold as $\mathcal{S}>0.5$ for selecting positive images in all three subtasks. As shown in Figure 6(b), the decrease in data quality leads to degraded T2I performance of both post-SFT and post-RL models. This highlights that the quality of SFT data is more important than its quantity.
@@ 5 Discussion
Recent work zhuang2025vargpt ; pan2024auto ; wang2025discrete ; deng2025emerging ; tong2024metamorph ; wu2024vila ; wu2024liquid has extensively discussed how to facilitate better unified visual comprehension and generation. However, the definition of “unified” seems to be reduced to whether an MLLM possesses both visual comprehension and generation capabilities simultaneously, with evaluation protocols typically designed for single tasks in isolation. We think it dilutes the true meaning of “unified”. The essence of unification should be reflected in the synergistic enhancement between the two capabilities, such as comprehension and generation improving each other through collaborative efforts.
This leads to a critical question: what kind of training paradigm can promote such unification? From Tables 4 (Rows 1-5) and 3, we observe that although we can endow models with the foundational ability to collaborate on visual comprehension and generation through SFT, the model appears to merely mechanically mimic and combine these two skills. The combination does not bring about any substantial improvement to either capability. Specifically, the performance of Janus-Pro-SFT-7B in both the visual comprehension task of image-text semantic consistency judgment and the text-to-image generation task shows very incremental improvement compared to Janus-Pro-7B. However, through RL, we encourage the model to spontaneously integrate the two capabilities to explore reasoning pathways, and merely provide incentives as appropriate guidance. We find that the model learns how to better coordinate these two capabilities, resulting in stronger text-to-image generation and image semantic understanding abilities.
Therefore, we argue that RL holds the potential to unlock genuine unified visual comprehension and generation. However, the current experiments are insufficient to conclusively validate this assertion. Owing to limited computational resources, our conclusion is based solely on simple image generation tasks and text-image semantic alignment evaluation tasks. Nevertheless, the experiments do provide some preliminary evidence for this conclusion. It is noteworthy that even the most advanced research deng2025emerging on unified visual comprehension and generation currently still focuses on supervised text-image alignment pretraining to emerge such unification. Given adequate computational resources, we think that large-scale RL could feasibly be employed to achieve a synergistic enhancement of visual comprehension and generation for genuine unification.
6 Conclusion and Future Work
@@
Future Work.
@@
  • Limitation in Section D.
@@ B.2 Reward Calculation
The overall design philosophy of our reward model is to leverage QA-based visual comprehension [6, 26, 36, 25, 41, 8, 31, 30] models, which will return a consistency score $\texttt{R}^{QA}(\cdot)\in[0,1]$ for each text-image pair, to assess the accuracy of the generated image and the MLLM’s self-evaluation. And we provide incentives for the final output images along with their preceding reasoning chains, incorporating bi-level reward scores: visual generation reward $\texttt{R}^{Gen}$ and visual comprehension reward $\texttt{R}^{Comp}$. Assume that we permit the MLLM to perform up to a maximum of $T$-round image generations and the MLLM determines that it has correctly generated the image in the $K$-th round (resulting in a total of $K$ images $\{\mathcal{I}^{i}\}_{i=1}^{K}$, $T\geq K$), we have: (1) The generation reward $\texttt{R}^{Gen}=\sum_{i=1}^{K-1}\texttt{R}^{QA}(\mathcal{I}^{i})+(T-K+1)\times\texttt{R}^{QA}(\mathcal{I}^{K})$. The $K$-th image is assigned a potentially larger weight because it represents the final output with higher importance. (2) The comprehension reward $\texttt{R}^{Comp}=\sum_{i=1}^{K}\big{(}\mathbf{1}-|\texttt{R}^{QA}(\mathcal{I}^{i})-\texttt{SE}(\mathcal{I}^{i})|\big{)}\times T/K$, where SE()=0or $\texttt{SE}(\mathcal{I})=0\ \text{or}\ 1$ is the self-assessment on whether the generated image is semantically aligned with the text.
@@
Hyper-parameters SFT for T2I RL for T2I SFT for Image Editing RL for Image Editing Optimizer AdamW AdamW AdamW AdamW Optimizer param. $\beta_{1}=0.9,\beta_{2}=0.95,\epsilon=1\mathrm{e}{-6}$ $\beta_{1}=0.9,\beta_{2}=0.95,\epsilon=1\mathrm{e}{-6}$ Peak LR 2.0e-5 6.0e-6 2.0e-5 1.0e-5 Convert LR - 2.0e-6 - 2.0e-6 Convert step - 400 - 300 Min LR 2.0e-7 2.0e-7 2.0e-7 2.0e-7 LR scheduler Cosine Linear+Cosine Cosine Linear+Cosine Batch size 128 128 128 128 Group size - 7 - 8 $\beta$ 0.05 - 0.05 Training Steps 50K 3K 40K 2.2K Warmup Steps 1000 100 1000 100 Weight decay 0.05 0.05 0.05 0.05 Gradient clipping 1.0 1.0 1.0 1.0 Numerical precision bfloat16 bfloat16 bfloat16 bfloat16 Resource Usage 8 NVIDIA A800 32 NVIDIA A800 8 NVIDIA A800 32 NVIDIA A800
@@ B.4 Inference
Janus-Pro-R1 can generate both text and images with two separate MLLM heads: an image head for predicting image tokens and a text head for predicting text tokens. We transform the text-to-image generation into an introspective process [40, 37]. During inference, when provided with a prompt, the model initially engages the image head to perform an autoregressive prediction of $576$ visual tokens to generate the first image. Subsequently, it switches to the text head to assess whether the generated image aligns with the semantic of the prompt and to predict a relevant reason. If the model determines that the image is consistent with the prompt, the inference is terminated, and the generated image is output. Conversely, if the model deems the image inconsistent, it switches back to the image head to predict another $576$ visual tokens for a new image generation attempt. The model then reverts to the text head to reassess the semantic alignment and predict a reason. This cycle of switching between the image and text heads continues iteratively. Of course, to prevent infinite image generation, the model is constrained to generate a maximum of $3$ images for each inference.
@@ C.2 Evaluation Benchmarks
Appendix D Limitation
Firstly, in the text-to-image generation task, our constructed prompt data contains relatively few instances related to counting. This directly results in our model’s counting metric on the Geneval dataset lagging behind SOTA models. To address this issue, we will further specifically construct a set of counting-related data to enhance the corresponding capabilities for counting-related prompts. Secondly, in the image editing task, although the filtered training datasets meet the requirements for image editing, many of the data samples have poor aesthetic appeal. This leads to some edited images lacking in aesthetic quality. To improve this, we plan to use GPT-4o to create a set of high-aesthetic image editing data in the future to enhance the natural appearance of the edited images.

### v2

# Janus-Pro-R1: Advancing Collaborative Visual Comprehension and Generation via Reinforcement Learning

@@ Abstract
11footnotetext:  Equal Contribution.33footnotetext:  Project Leader.22footnotetext:  Corresponding Author.
@@ 2 Related Work
Recent studies focus on unifying visual comprehension and generation within a unified MLLM. Some efforts ge2024seed ; sun2024generative ; tong2024metamorph cascade external diffusion models after the output of MLLMs for visual generation, while some approaches chen2025janus ; pan2025generative ; zhuang2025vargpt ; wang2024emu3 ; pan2025focusdiff tokenize images into a discrete space and then conducte unified AR for text and vision. Others xie2024show ; zhou2024transfusion try to integrate the objective of AR and diffusion into a single model. However, visual comprehension and generation still remain largely independent without collaborative synergy, and the strong reasoning mechanism of LLMs fails to revolutionize visual generation. In this paper, we aim to promote collaboration between these two capabilities, enabling a genuine CoT for visual generation and unlocking its Aha moments.
@@ 3.1 Supervised Fine-Tuning
Task-II: Self-evaluation of Text-Image Consistency: We aim to enhance the capability of MLLM to determine whether given images are semantically consistent with the text, also providing the rationale for its judgment. To construct training data, for each prompt we first select (t1)(t-1) images $\{\mathcal{I}^{j},\mathcal{R}^{j}\}_{j=1}^{t-1}$ with $\mathcal{S}<0.5$ as the preceding context ($t\geq 1$). Subsequently, we randomly select one positive image ($\mathcal{S}\geq 0.7$) and one negative image ($\mathcal{S}<0.5$), respectively, as the target image $\mathcal{I}_{t}$ to be evaluated. The objective function is $p(y^{\mathcal{T}^{t}})=\sum_{j=1}^{S^{\mathcal{T}^{t}}}\log P_{\theta}(y^{\mathcal{T}^{t}}_{j}|y^{\mathcal{T}^{t}}_{<j},\mathcal{P},\{\mathcal{I}^{k},\texttt{No},\mathcal{R}^{k}\}_{k=1}^{t-1},\mathcal{I}^{t})$, where $\mathcal{T}^{t}$ is $(\text{Yes},\mathcal{R}^{t})$ for positive images and $(\text{No},\mathcal{R}^{t})$ for negative images.
@@ 3.2 Reinforcement Learning
(2) The comprehension reward $\texttt{R}^{Comp}=\sum_{i=1}^{K}\big(\mathbf{1}-|\texttt{R}^{QA}(\mathcal{I}^{i})-\texttt{SE}(\mathcal{I}^{i})|\big)\times T/K$, where $\texttt{SE}(\mathcal{I})=0\ \text{or}\ 1$ is the self-assessment on whether the generated image is semantically aligned with the text.
@@
[…]
@@ 4 Experiments
We employ Janus-Pro-7B as the backbone, developing Janus-Pro-R1 and Janus-Pro-R1-Edit, excelling in text-to-image generation (§ 4.1), image editing (§ 4.2), and image semantic evaluation (§ 4.3). More details are given in Appendix C and  E.
@@ 4.1 Introspective Text-to-Image Generation
We first conduct an automated metric evaluation on 3 text-to-image benchmarks: GenEval ghosh2023geneval , T2I-CompBench huang2023t2i , and DPG-Bench hu2024ella . The comparison results against both diffusion-based and MLLM-based methods, as well as methods incorporating CoT into image generation, are presented in Table 1. We have the following observations:
@@
In Figure 3 and Figure 7, we present qualitative examples of Janus-Pro-R1 to trigger Aha moments within its reasoning chains to generate superior images. The model could leverage its visual comprehension capabilities to accurately identify the issues in its initial-generated images, then unleash the visual generation capabilities to output a more accurate image. Even if the newly generated image still fails to meet the requirements, the model can trigger a second Aha moment, re-evaluating the issues and repeating the visual generation to produce a fully compliant image.
@@
Method
T2I
Model
Distance $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ MSE $\downarrow$ SSIM $\uparrow$ Whole $\uparrow$ Edited $\uparrow$
@@ 4.2 Image Editing
Figure 4: Our model achieve a stable trade-off between fidelity and editability in multi-turn editing.
@@ 4.3 Janus-Pro-R1 as a Stronger Image Semantic Evaluator
Furthermore, we collect several hundred pairs of semantically mismatched text-image pairs. We require the above four models to determine whether the semantics of the image and text are matched and to provide relevant reasons. We then use GPT-4o as the evaluation model to score the reasons provided by the four models from 0 (unreasonable) to $1$ (reasonable). As shown in Table 3, Janus-Pro-R1 outperforms both Janus-Pro-7B and Janus-Pro-SFT in providing reasons that are deemed more reliable by GPT-4o. It even approaches the performance of the comprehension-only MLLM, i.e., InternVL2.5-8B. These results indicate that (1) compared to the backbone model and even a comprehension-only MLLM, Janus-Pro-R1 more closely aligns with the standard metrics within GenEval, achieving greater evaluation precision; and (2) the visual comprehension capability of our model is significantly enhanced after reinforcement learning. In contrast, SFT only endows the MLLM with the basic capability for combining visual comprehension and generation, resulting in only incremental improvements to its inherent multimodal understanding ability.
@@ 4.4 In-Depth Analysis
During SFT, when selecting positive images for subtask training, we establish a higher threshold larger than 0.5 of the semantic consistent score ($\mathcal{S}\geq 0.7$ for comprehension tasks and $\mathcal{S}\geq 0.8$ for generation tasks). Though a higher threshold reduces the quantity of available training data, we find that it is crucial for improving performance. In contrast, we also set the threshold as $\mathcal{S}>0.5$ for selecting positive images in all three subtasks. As shown in Figure 5(b), the decrease in data quality leads to degraded T2I performance of both post-SFT and post-RL models. This highlights that the quality of SFT data is more important than its quantity.
@@
The essence of unification for comprehension and generation should be reflected in the synergistic enhancement between the two capabilities. So what kind of training paradigm can promote such unification? From Tables 4 (Rows 1-5) and 3, we observe that although we can endow models with the foundational ability to collaborate on visual comprehension and generation through SFT, the model appears to merely mechanically mimic and combine these two skills. The combination does not bring about any substantial improvement to either capability. Specifically, the performance of Janus-Pro-SFT-7B in both the visual comprehension task of image-text semantic consistency judgment and the text-to-image generation task shows very incremental improvement compared to Janus-Pro-7B. However, through RL, we encourage the model to spontaneously integrate the two capabilities to explore reasoning pathways, and merely provide incentives as appropriate guidance. We find that the model learns how to better coordinate these two capabilities, resulting in stronger text-to-image generation and image semantic understanding abilities.
Therefore, we argue that RL holds the potential to unlock genuine unified visual comprehension and generation. However, the current experiments are insufficient to conclusively validate this assertion. Owing to limited computational resources, our conclusion is based solely on simple image generation tasks and text-image semantic alignment evaluation tasks. Nevertheless, the experiments do provide some preliminary evidence for this conclusion. Given adequate computational resources, we think that large-scale RL could feasibly be employed to achieve a synergistic enhancement of visual comprehension and generation for genuine unification, empowering MLLMs to demonstrate powerful image understanding and generation capabilities comparable to GPT-4o.
5 Conclusion
@@
Acknowledgements
This work was supported by the Key R&D Projects in Zhejiang Province (No. 2024C01106, 2025C01030), the NSFC (62272411), and the Zhejiang NSF (LRG25F020001).
@@ Acknowledgements
  • More Experimental Results in Section D.
  • Limitation, Future Work, Broader Impacts and Safeguards in Section E.
@@ B.2 Reward Calculation
The overall design philosophy of our reward model is to leverage QA-based visual comprehension [6, 25, 36, 24, 42, 8, 30, 29] models, which will return a consistency score $\texttt{R}^{QA}(\cdot)\in[0,1]$ for each text-image pair, to assess the accuracy of the generated image and the MLLM’s self-evaluation. And we provide incentives for the final output images along with their preceding reasoning chains, incorporating bi-level reward scores: visual generation reward $\texttt{R}^{Gen}$ and visual comprehension reward $\texttt{R}^{Comp}$. Assume that we permit the MLLM to perform up to a maximum of $T$-round image generations and the MLLM determines that it has correctly generated the image in the $K$-th round (resulting in a total of $K$ images $\{\mathcal{I}^{i}\}_{i=1}^{K}$, $T\geq K$), we have: (1) The generation reward $\texttt{R}^{Gen}=\sum_{i=1}^{K-1}\texttt{R}^{QA}(\mathcal{I}^{i})+(T-K+1)\times\texttt{R}^{QA}(\mathcal{I}^{K})$. The $K$-th image is assigned a potentially larger weight because it represents the final output with higher importance. (2) The comprehension reward $\texttt{R}^{Comp}=\sum_{i=1}^{K}\big(\mathbf{1}-|\texttt{R}^{QA}(\mathcal{I}^{i})-\texttt{SE}(\mathcal{I}^{i})|\big)\times T/K$, where $\texttt{SE}(\mathcal{I})=0\ \text{or}\ 1$ is the self-assessment on whether the generated image is semantically aligned with the text.
@@
Hyper-parameters SFT for T2I RL for T2I SFT for Image Editing RL for Image Editing Optimizer AdamW AdamW AdamW AdamW Optimizer param. $\beta_{1}=0.9,\beta_{2}=0.95,\epsilon=1\mathrm{e}{-6}$ $\beta_{1}=0.9,\beta_{2}=0.95,\epsilon=1\mathrm{e}{-6}$ Peak LR 2.0e-5 6.0e-6 2.0e-5 1.0e-5 Convert LR - 2.0e-6 - 2.0e-6 Convert step - 400 - 300 Min LR 2.0e-7 2.0e-7 2.0e-7 2.0e-7 LR scheduler Cosine Linear+Cosine Cosine Linear+Cosine Batch size 128 128 128 128 Group size - 7 - 8 $\beta$ - 0.05 - 0.05 Training Steps 50K 3K 40K 2.2K Warmup Steps 1000 100 1000 100 Weight decay 0.05 0.05 0.05 0.05 Gradient clipping 1.0 1.0 1.0 1.0 Numerical precision bfloat16 bfloat16 bfloat16 bfloat16 Resource Usage 8 NVIDIA A800 32 NVIDIA A800 8 NVIDIA A800 32 NVIDIA A800
@@ B.4 Inference
Janus-Pro-R1 can generate both text and images with two separate MLLM heads: an image head for predicting image tokens and a text head for predicting text tokens. We transform the text-to-image generation into an introspective process [41, 37, 38]. During inference, when provided with a prompt, the model initially engages the image head to perform an autoregressive prediction of $576$ visual tokens to generate the first image. Subsequently, it switches to the text head to assess whether the generated image aligns with the semantic of the prompt and to predict a relevant reason. If the model determines that the image is consistent with the prompt, the inference is terminated, and the generated image is output. Conversely, if the model deems the image inconsistent, it switches back to the image head to predict another $576$ visual tokens for a new image generation attempt. The model then reverts to the text head to reassess the semantic alignment and predict a reason. This cycle of switching between the image and text heads continues iteratively. Of course, to prevent infinite image generation, the model is constrained to generate a maximum of $3$ images for each inference.
@@ C.2 Evaluation Benchmarks
Figure 7: More qualitative examples of introspective text-to-image generation that triggers Aha moments.
Appendix D More Experimental Results
In Figure 7, we present more qualitative examples of Janus-Pro-R1 to trigger Aha moments within its reasoning chains to generate superior images. Our model could leverage its visual comprehension capabilities to accurately identify the issues in its initial-generated images, then unleash the visual generation capabilities to output a more accurate image.
Figure 8 presents a direct qualitative comparison between Janus-Pro-7B and our Janus-Pro-R1-7B on the text-to-image generation task, with both short and long captions. It can be observed that compared to Janus-Pro-7B, after unlocking the Aha moments with CoT via a two-stage training paradigm, our model not only generates images that are more semantically aligned with the text but also achieves higher aesthetic quality.
Appendix E Limitation, Future Work, Broader Impacts and Safeguards
Limitations.
Firstly, in the text-to-image generation task, our constructed prompt data contains relatively few instances related to counting. This directly results in our model’s counting metric on the Geneval dataset lagging behind SOTA models. To address this issue, we will further specifically construct a set of counting-related data to enhance the corresponding capabilities for counting-related prompts. Secondly, in the image editing task, although the existing datasets meet the requirements for image editing, many of the data samples have poor aesthetic appeal. This leads to some edited images lacking in aesthetic quality. To improve this, we plan to use GPT-4o to create a set of high-aesthetic image editing data in the future to enhance the natural appearance of the edited images.
Future Work.
Broader Impacts.
This study does not raise any ethical concerns. The research does not involve subjective assessments or the use of private data. Only publicly available datasets are utilized for experimentation.
Safeguards.
A major societal concern with this technology lies in its potential for misuse, particularly in fabricating unauthorized images that could lead to misinformation, privacy breaches, and other damaging consequences. To counter these threats, it is crucial to develop strong ethical standards and implement ongoing surveillance.